Your US-based firm is considering investing in a project run by its Canadian subsidiary. This project will cost CAD 26M to set up today and will pay out CAD 29M in one year. This project will be all equity financed, with the parent firm taking a 70% equity stake, and the Canadian subsidiary will be taking a 30% equity stake. The spot rate is currently USD 0.77 per CAD, and you expect that it will be USD 0.84 per CAD in one year. Your USD discount rate for projects in First World foreign countries is 14%, and your Canadian subsidiary’s discount rate for domestic projects is 13%

c.You believe that the payout for this project could be CAD 1M higher or lower than expected, and that the exchange rate one year from today could be USD 0.12 per CAD higher or lower than expected. If you are concerned about reducing the chance of making a negative NPV investment for the parent firm, should you concentrate on eliminating exchange rate risk or boosting sales?
Answer #1

The question requires evaluating NPV for the parent company. Thus, where foreign cash-flows are involved and NPV to be calculated using the parent company in US, the following steps are to be done:

1. Convert CAD cash-flows to USD using the applicable exchange rate

2. Discount the USD cash-flows (as per step 1) using the US interest rate to arrive at the NPV.

A B 1 26,000,000 29,000,000 70% 30% 0.77 0.84 2 Project Cost (CAD) 3 Pay-out in 1 year 4 Equity stake of US Parent 5 Equity s

Net Present Value with the current investment and cash-inflow at the given spot rate and expected spot rate is USD 943,895.

Workings:

A B 1 2 Project Cost (CAD) =26*10^6 3 Pay-out in 1 year =29*10^6 4 Equity stake of US Parent 0.7 5 Equity stake of Canada Sub

The payout for this project could be CAD 1M higher or lower than expected, and that the exchange rate one year from today could be USD 0.12 per CAD higher or lower than expected. If you are concerned about reducing the chance of making a negative NPV investment for the parent firm, should you concentrate on eliminating exchange rate risk or boosting sales?

Lets find the decrease in NPV if there is CAD 1M drop in sales (one year pay out) with the same expected spot rate of 0.84 and also same one year pay-out of $29 M with drop in expected spot rate by 0.12 :

A B с 1 2 Project Cost (CAD) 3 Pay-out in 1 year 4 Equity stake of US Parent 5 Equity stake of Canada Subsidiary 6 Spot rate

NPV falls by $515,789 if the sales decrease by CAD 1 million and it falls by $2,136,842 if the exchange rate falls by 0.12 (to USD 0.74 per CAD),.

The decrease in NPV is significantly higher if the exchange rate falls by USD 0.12 per CAD than the reduction in sales. The USD 0.12 reduction causes the resultant NPV to go negative at $1,192,947 ($943,895-$2,136,842). Thus, the concentration should be on eliminating exchange rate risk than boosting sales.

Workings:

A B C 1 2 Project Cost (CAD) =26*10^6 3 Pay-out in 1 year =29*10^6 4 Equity stake of US Parent 0.7 5 Equity stake of Canada S